# MambaXCTrack: Mamba-based Tracker with SSM Cross-correlation and Motion Prompt for Ultrasound Needle Tracking

### v1

@@ III Experiments and Results
III-A Experiment Setup and Dataset Collection
Figure 3: Experiment setup. The in-plane case (imaging plane aligned parallel with needle trajectory) is shown as an example.
@@ III-A Experiment Setup and Dataset Collection
TABLE II: Average errors and standard deviations in mm.
Method Phantom Tissue Mean
SiamRPN++ [10] 4.11$\pm$3.05 3.21$\pm$2.58 3.70$\pm$2.83
SiamCAR [17] 2.26$\pm$1.55 2.23$\pm$1.37 2.24$\pm$1.47
SiamBAN [11] 0.70$\pm$0.54 2.09$\pm$1.55 1.33$\pm$1.01
SiamAttn [18] 0.63$\pm$0.67 1.62$\pm$1.33 1.08$\pm$0.98
Yan et.al. [6] 0.98$\pm$0.94 2.83$\pm$2.21 1.84$\pm$1.53
STMTrack [29] 0.41$\pm$0.27 2.32$\pm$1.52 1.28$\pm$0.86
SwinTrack [30] 0.37$\pm$0.17 1.54$\pm$1.53 0.90$\pm$0.81
MixFormerV2 [12] 0.26$\pm$0.25 1.37$\pm$1.51 0.76$\pm$0.84
MambaXCTrack 0.22$\pm$0.14 0.49$\pm$1.01 0.34$\pm$0.55
@@
The experiment setup is shown in Fig. 3. The US imaging was performed with a Verasonics Vantage 32 LE US machine with a Mindray C5-2 US probe. An electromagnetic (EM) localization system (Aurora, NDI Inc.) was adopted to collect the needle tip’s position as the ground truth, using an 18-gauge needle with a 5-DoF EM tracker embedded on its tip. The needle was attached to a linear manipulator for motorized insertion. The US probe was installed on a 3-axis manipulator to move the probe for needle visualization when needed. The phantom was made of agar of 3% mass concentration and the tissue is a slice of fresh pork. Silica powder was added to agar to simulate speckles in real anatomical tissue.
@@
During the experiment, the data of phantom-only experiment and tissue experiment were collected separately regarding three cases, i.e., in-plane (static), in-plane (moving), and out-of-plane [19]. For cases of in-plane (moving) and out-of-plane, the probe moved at the same speed as the horizontal velocity of needle insertion. For in-plane (static), only the needle moved. The needle insertion was performed with three angles ($0^{\circ}$, $30^{\circ}$, and $60^{\circ}$) and three velocities ($0.4$, $1$, and $2$ mm/s) for each case. For each pair of angle and velocity, at least 6 procedures were conducted. Each needle insertion length is more than 120 mm to ensure a thorough procedure. 108 procedures have been done in total, including 51 in tissue and 57 in phantom, which results in 108 videos ($492\times 856$, 30 FPS) with 87330 frames. The ground truth of the needle tip position was acquired by the EM tracking system, which has been verified to have a satisfactory RMSE of 0.76 mm in our environment.
@@
To train the model, the videos were first sampled by 10 to get 8733 images to remove redundancy. The dataset was then split into training, validation, and testing sets in the ratio 7:1:2. A video/setup-level data splitting was conducted to ensure images from the same video are not divided into different sets, and videos with different experiment setups are evenly divided into different sets. The program was implemented with PyTorch. All the models involved in this paper were trained on four NVIDIA RTX 4090 GPUs with the same training strategy (350 epochs, batch size 32, AdamW optimizer). The learning rate was set to 3e-4 with a backbone learning rate of 3e-5, which dropped by 10 after 200 epochs. Scaling, blur, and position shifting were adopted as image augmentation. Gaussian noise was added to the historical motion sequence.
@@ III-B Results
The evaluation was conducted based on three metrics that are commonly adopted in tracking evaluation, namely area under curve (AUC) [31], precision (P) [32], and normalized precision ($\textit{P}_{\textit{norm}}$) [32]. AUC reports the needle tracking success rate by evaluating the area under the Success Plot curve as defined in [31], which offers a comprehensive evaluation of tracking robustness. Higher AUC suggests that the tracker keeps the needle tip tracked for a larger portion of the frames. Beyond success rate by AUC, tracking error is evaluated by P and $\textit{P}_{\textit{norm}}$ [32]. P evaluates the accuracy of predicted position measured by Center Location Error (CLE), which is the Euclidean distance between the predicted and ground truth position. P is generated by calculating the percentage of frames where the CLE is less than a specific threshold, which is set to a conventional value of 20 pixels. $\textit{P}_{\textit{norm}}$ adjusts P to account for the scale of the object by normalizing CLE with object size, providing a more consistent measure across varying object sizes. Higher P and $\textit{P}_{\textit{norm}}$ can be interpreted as needle tip tracking with higher accuracy. Besides, the average tracking errors and standard deviations in millimeter were also reported to provide a better comparison in terms of physical metrics. Several state-of-the-art natural object trackers were used for comparison, including state-of-the-art transformer-based tracker MixFormerV2 [12], tracker with motion prompts (SwinTrack [30]), tracker with aggregated historical templates (STMTrack [29]), classical Siamese tracker SiamRPN++ [10] and its improved variants [11, 17, 18]. The state-of-the-art tracker in US-based needle tracking [6] was also used as a comparison.
@@
As the results show in Tab. I and Tab. II, the proposed MambaXCTrack achieves SOTA performance in almost all metrics of all evaluations. In the phantom environment with less background noise and more stable needle appearance, while most methods offer satisfactory tracking performance in the out-of-plane cases, a few trackers, including MambaXCTrack, distinguish themselves with superior performance in the in-plane cases. SwinTrack and MixFormerV2 were ranked second in the In-plane (static) and In-plane (moving) cases, respectively, while MambaXCTrack outperforms both to rank first in both cases. This achievement can be attributed to the higher modeling efficiency of the Mamba-based structure compared with the transformer architecture adopted in both SwinTrack and MixFormerV2. MambaXCTrack also achieves a 0.22$\pm$0.14 mm tracking error and standard deviation (SD). This means that the estimated tip position by MambaXCTrack can be maintained within the 18-gauge needle tip (with a cross-sectional radius of 0.625 mm) in almost all frames.
@@
TABLE III: Ablation studies of the baseline model against seven variations.
@@
In the tissue environment that is much more challenging than the phantom, MambaXCTrack consistently achieves the best and second-best performance in nearly all evaluations. MambaXCTrack significantly outperforms Yan et.al. [6] by 125.5%, 104.8%, and 68.1% in AUC for all three cases. While both MambaXCTrack and the work in [6] integrate motion information to address intermittent invisibility issue, the superior performance by MambaXCTrack is mainly attributed to longer modeling distance of SSMX-Corr than transformer in [6] and better generalizability of implicit low-level motion descriptor than explicit motion predictor in [6]. As shown in Tab. II, MambaXCTrack achieves a tracking error of 0.49$\pm$1.01 mm in tissue experiments, outperforming the second-best MixFormerV2 by 64.2% and 33.1%. The offset of the predicted position relative to the tip ground truth is kept within a distance of the needle cross-sectional radius, better than MixFormerV2 which has an average error of 2.2 times the needle radius.
@@
Figure 4: Tracking demonstration of MambaXCTrack against three SOTA trackers in three scenarios. The view is zoomed in. The upper row is the forward procedure, while the bottom is the backward. The ground truth of tip position and needle trajectory are marked. See more examples in supplementary video.
A tracking demonstration is shown in Fig. 4 including three examples [^1] More tracking demonstration is provided in the supplementary video. that demonstrate different challenges in US needle tracking. The first example of In-plane (static) / 0 / Phantom in the first column shows a scenario with interference from needle traces. MambaXCTrack overcomes the interference from needle traces, which arise from the hollow regions created by needle retraction and have a very similar appearance to the needle tip. This interference has led to STMTrack’s failures, whereas our tracker maintains successful tracking. For procedures done within the tissue, the distraction from noise and artifacts becomes more severe, and the needle visibility is degraded. The second example of In-plane (moving) / 30 / Tissue demonstrates this distraction and degradation, where the background is filled with noise from high-intensity speckles caused by the density variation of tissue. The backward procedure of this example also shows the intermittent invisible needle tip, since the imaging plane and needle tend to be misaligned in this dynamic case, where both probe and needle are moving. MambaXCTrack is the only one among the four that maintains successful tracking even when the tip is intermittently invisible. A similar case of tracking intermittently disappearing needle is also demonstrated in backward of out-of-plane / 60° / tissue, where MixFormerV2 lost tracking but MambaXCTrack was able to keep the target tracked. When inferencing, MambaXCTrack runs at 34.9 FPS on average, satisfying the real-time tracking requirement.
@@ III-C Ablation Study
Extensive ablation studies are conducted regarding the baseline model against seven variations $v_{1}$ to $v_{7}$, with their configurations and results shown in Tab. III. The first three are adopted to test the proposed SSMX-Corr and CIS. CIS is kept in $v_{1}$ but $m$ is concatenated between $z$ and $x_{i}$, different from the baseline ($[m,z,x_{i}]$), to evaluate different orders of information fusion. CIS is removed in $v_{2}$ and three feature maps are simply concatenated together into a sequence $[m,z,x]$. CIS is also removed in $v_{3}$ and SSMX-Corr is replaced by a convolutional X-Corr [10]. The results of these variations show the effectiveness of CIS and SSMX-Corr. Performance degradation in $v_{1}$ demonstrates that the tracking benefits more from motion information by concatenating $m$ at the beginning of $z$ and $x_{i}$. With CIS removed ($v_{2}$), performance degradation is observed in almost all evaluations, showing that tracking can be improved by building pixel-wise interaction between $z$ and $x$. Performance drop is also observed in $v_{3}$, especially for the tissue experiments, where noise distraction is much more severe. It demonstrates that tracking under challenging environments is greatly improved by learning potential distant visual cues utilizing SSMX-Corr’s long-range modeling capability.
Ablation regarding the proposed low-level motion descriptor is also carried out, namely $v_{4}$ to $v_{7}$. The motion descriptor is removed in $v_{4}$. $v_{5}$ keeps the motion descriptor but uses raw motion without preprocessing it like the baseline. Motion descriptors that cover different time durations are tested in $v_{6}$ (longer time, $T=120$) and $v_{7}$ (shorter time, $T=30$). A considerable degradation is reported in $v_{4}$, demonstrating that tracking is greatly enhanced by learning from historical low-level motion descriptors. Accuracy drops are also reported in almost all metrics of $v_{5}$, showing the improvement brought by extracting low-level descriptors from raw motion. The results of $v_{6}$ and $v_{7}$ demonstrate the length of baseline ($T=60$) is a more suitable time span for motion descriptors in US needle tracking which covers 2 seconds. This is likely because moderate-duration motion provides sufficient cues while avoiding interference from long-term tracking noise.
@@ III-D Discussion
MambaXCTrack outperforms a series of SOTA trackers including CNN-based trackers and transformer-based trackers. Its superior performance can be attributed to two main factors. Firstly, the SSMX-Corr with CIS effectively addresses the unique domain characteristics of US images, i.e., the local feature from the target can become unpredictably unreliable, searching for semantic cues over long distances significantly enhances tracking accuracy. Inspiration for long-range searching is straightforward and intuitive, which directly comes from the trackers that utilize needle shaft segmentation for guidance [15, 16]. These studies suggest that while the needle tip may occasionally become obscured, the larger needle shaft remains partially visible, providing essential directional axial information for tracking. Instead of direct segmentation, which faces challenges from vulnerability to noise and limited training data, SSM is employed to endow the model with long-range modeling capabilities, allowing it to learn from distant semantic features including the needle shaft. The existing ConvX-Corr adopted in SiamRPN++ [10] is a local pixel-wise operation that convolves $x$ with kernel $z$. While it is considered effective for general tasks like real-world object tracking, its limited search range proves inadequate for US needle tracking, where images are often degraded by noise and artifacts, rendering local fine features unreliable. This disadvantage can be observed from the performance of those four ConvX-Corr-based trackers ([10, 17, 11, 18]) and the ablation of ConvX-Corr ($v_{3}$). By utilizing SSMX-Corr, MambaXCTrack overcomes these challenges by identifying useful long-range semantic cues. With the further integration of CIS, SSMX-Corr retains the advantages of ConvX-Corr, gaining positional inductive bias from this convolution-like operation.
@@
The second aspect is the implicit low-level motion descriptor. When the needle temporarily disappears, historical motion becomes more critical than visual information. This intermittent visibility can result from obstructions by anatomical structures or deviations in US imaging plane. In such cases, relying on motion information rather than vision is a more effective strategy. Ablation studies demonstrate that integrating historical low-level motion descriptors with an appropriate time duration enhances tracking. SwinTrack [30] also incorporates motion information yet leads to worse performance than ours. This is partly because SwinTrack utilizes raw motion without first extracting low-level descriptors to mitigate disturbance from absolute value, which can harm generalization. Additionally, its information fusion framework is less efficient than the SSM used in our method, which is inherently suited for long-sequence modeling. STMTrack [29] adopts a template bank to store the historical frames, focusing on historical visual rather than motion information. However, it performs even worse than SwinTrack [30]. This underperformance likely arises from error accumulation due to aggregating historical US images, which contains severe noise and artifacts. Yet historical motion suffers less from error accumulation thanks to its simpler distribution and characteristic that does not contain high-level information from visual features. In challenging environments, leveraging simpler historical information is more effective than integrating high-level features. However, while the tracker proposed by Yan et al. [6] also utilizes motion information for ultrasound needle tracking, its performance is unsatisfactory, excelling only in a limited number of metrics. This suboptimal performance is likely attributed to its explicit motion prediction structure, which prevents the model from being fully end-to-end, resulting in error propagation and system instability. Directly predict future motion can also cause degraded generalizability when encountering videos from unseen domain, which contain motion that hard to be adapted by a simple motion predictor. On the contrary, MambaXCTrack leverages implicit motion integration to ensure an end-to-end structure with better stability and generalization by facilitating smoother information fusion through its use of SSM.

### v2

@@ III Experiments and Results
III-A Experimental Setup and Dataset Collection
Figure 3: Experimental setup. The in-plane case (imaging plane aligned parallel with needle trajectory) is shown as an example.
@@ III-A Experimental Setup and Dataset Collection
TABLE II: Average errors and standard deviations (mm) of the tracking, which are computed over the testing set with 1747 samples. Inference speed is reported in FPS.
Method Phantom Tissue Mean FPS
SiamRPN++ [10] 4.11$\pm$3.05 3.21$\pm$2.58 3.70$\pm$2.83 27.9
SiamCAR [20] 2.26$\pm$1.55 2.23$\pm$1.37 2.24$\pm$1.47 48.3
SiamBAN [19] 0.70$\pm$0.54 2.09$\pm$1.55 1.33$\pm$1.01 30.7
SiamAttn [21] 0.63$\pm$0.67 1.62$\pm$1.33 1.08$\pm$0.98 33.2
Yan et.al. [6] 0.98$\pm$0.94 2.83$\pm$2.21 1.84$\pm$1.53 24.0
STMTrack [33] 0.41$\pm$0.27 2.32$\pm$1.52 1.28$\pm$0.86 22.1
SwinTrack [34] 0.37$\pm$0.17 1.54$\pm$1.53 0.90$\pm$0.81 59.5
MixFormerV2 [11] 0.26$\pm$0.25 1.37$\pm$1.51 0.76$\pm$0.84 81.0
MambaXCTrack 0.22$\pm$0.14 0.49$\pm$1.01 0.34$\pm$0.55 34.9
@@
The experimental setup is shown in Fig. 3. Ultrasound imaging was performed with a Verasonics Vantage 32 LE US machine with a Mindray C5-2 US probe. This probe has a transmit frequency of 3.5 MHz with 96 transducer elements. During imaging, a plane wave sequence with an imaging depth of 150 mm was utilized, and the sampling frequency was set to 25 MHz. An electromagnetic (EM) localization system (Aurora, NDI Inc.) was adopted to collect the needle tip’s position as the ground truth, using an 18-gauge needle with a 5-DoF EM tracker embedded on its tip. The needle was attached to a linear manipulator for motorized insertion. The US probe was installed on a 3-axis manipulator to move the probe for needle visualization when needed. The phantom was made of agar of 3% mass concentration and the tissue was a slice of fresh pork. Silica powder was added to agar to simulate speckles in real anatomical tissue.
@@
During the experiment, the data of the phantom-only experiment and tissue experiment were collected separately regarding three cases, i.e., in-plane (static), in-plane (moving), and out-of-plane [22]. For cases of in-plane (moving) and out-of-plane, the probe moved at the same speed as the horizontal velocity of needle insertion. For in-plane (static), only the needle moved. The needle insertion was performed with three angles ($0^{\circ}$, $30^{\circ}$, and $60^{\circ}$) and three velocities ($0.4$, $1$, and $2$ mm/s) for each case. For each pair of angle and velocity, at least 6 procedures were conducted. Each needle insertion length is more than 120 mm to ensure a thorough procedure. 108 procedures have been done in total, including 51 in tissue and 57 in phantom, which results in 108 videos ($492\times 856$, 30 FPS) with 87330 frames. The pixel resolution of each frame ($492\times 856$) corresponds to a field-of-view of 150 mm $\times$ 260 mm. The ground truth of the needle tip position was acquired by the EM tracking system, which has been verified to have a satisfactory RMSE of 0.76 mm in our environment.
@@
To train the model, the videos were first sampled by 10 to get 8733 images to remove redundancy. The dataset was then split into training (6113 samples), validation (873 samples), and testing (1747 samples) sets in the ratio 7:1:2. A video/setup-level data splitting was conducted to ensure images from the same video were not divided into different sets, and videos with different experiment setups were evenly divided into different sets. The program was implemented with PyTorch. All model training and inferencing was performed on a server running Ubuntu 22.04 with two Intel Xeon Platinum 8375C CPUs and four NVIDIA RTX 4090 GPUs installed. Inferencing only involves one GPU. All models were trained with the same training strategy (350 epochs, batch size 32, AdamW optimizer), and they used the same input resolution of $384\times 384$ for search images and $192\times 192$ for template images. The learning rate was set to 3e-4 with a backbone learning rate of 3e-5, where both of them were dropped by a factor of 10 after 200 epochs. The backbone is trained together with the network but adopts a smaller learning rate. Scaling, blur, and position shifting were adopted as image augmentation. Gaussian noise was added to the historical motion sequence. The augmentations were performed dynamically during the training process.
@@ III-B Results
The evaluation was conducted based on three metrics that are commonly adopted in tracking evaluation, namely area under curve (AUC) [35], precision (P) [36], and normalized precision ($\textit{P}_{\textit{norm}}$) [36]. AUC reports the needle tracking success rate by evaluating the area under the Success Plot curve as defined in [35], which offers a comprehensive evaluation of tracking robustness. Higher AUC suggests that the tracker keeps the needle tip tracked for a larger portion of the frames. Beyond success rate by AUC, tracking error is evaluated by P and $\textit{P}_{\textit{norm}}$ [36]. P evaluates the accuracy of predicted position measured by Center Location Error (CLE), which is the Euclidean distance between the predicted and ground truth positions. P is generated by calculating the percentage of frames where the CLE is less than a specific threshold, which is set to a conventional value of 20 pixels. $\textit{P}_{\textit{norm}}$ adjusts P to account for the scale of the object by normalizing CLE with the size of the ground-truth bounding box, providing a more consistent measure across varying object sizes. The ground-truth bounding box is defined based on the needle tip region that is visible in the ultrasound imaging. Higher P and $\textit{P}_{\textit{norm}}$ can be interpreted as needle tip tracking with higher accuracy. In addition to these three metrics, the average tracking errors and standard deviations in millimeters were also reported to provide a better comparison in terms of physical metrics. Several state-of-the-art natural object trackers were retrained on the proposed dataset and used for comparison, including state-of-the-art transformer-based tracker MixFormerV2 [11], tracker with motion prompts (SwinTrack [34]), tracker with aggregated historical templates (STMTrack [33]), classical Siamese tracker SiamRPN++ [10] and its improved variants [19, 20, 21]. A US-based needle tracker by Yan et al. [6] was also used as a comparison.
@@
As the results show in Tab. I and Tab. II, the proposed MambaXCTrack achieves SOTA performance in almost all metrics of all evaluations while maintaining satisfactory inference speed. In the phantom environment with less background noise and more stable needle appearance, while most methods offer satisfactory tracking performance in the out-of-plane cases, a few trackers, including MambaXCTrack, distinguish themselves with superior performance in the in-plane cases. SwinTrack and MixFormerV2 were ranked second in the in-plane (static) and in-plane (moving) cases, respectively, while MambaXCTrack outperforms both to rank first in both cases. This achievement can be attributed to the higher modeling efficiency of the Mamba-based structure compared with the transformer architecture adopted in both SwinTrack and MixFormerV2. MambaXCTrack also achieves a 0.22$\pm$0.14 mm tracking error and standard deviation (SD). This means that the estimated tip position can be maintained within the 18-gauge needle tip (with a cross-sectional radius of 0.625 mm) in almost all frames.
@@
TABLE III: Ablation studies of the baseline model (MambaXCTrack) against seven variations. The results of the phantom and tissue experiments are reported as mean values of the three insertion techniques.
@@
In the tissue environment that is much more challenging than the phantom, MambaXCTrack consistently achieves the best and second-best performance in nearly all evaluations. MambaXCTrack significantly outperforms Yan et.al. [6] by 125.5%, 104.8%, and 68.1% in AUC for all three cases. While both MambaXCTrack and the work in [6] integrate motion information to address the intermittent invisibility issues, the superior performance by MambaXCTrack is mainly attributed to longer modeling distance of SSMX-Corr than the transformer in [6] and better generalizability of an implicit low-level motion descriptor than an explicit motion predictor in [6]. As shown in Tab. II, MambaXCTrack achieves a tracking error of 0.49$\pm$1.01 mm in tissue experiments, outperforming the second-best MixFormerV2 by 64.2% and 33.1%. The offset of the predicted position relative to the tip ground truth is kept within a distance of the needle cross-sectional radius, better than MixFormerV2, which has an average error of 2.2 times the needle radius.
@@
Figure 4: Tracking demonstration of MambaXCTrack against three SOTA trackers in three scenarios. The view is zoomed in. The left column is the forward procedure, while the right is backward. The ground-truth tip position and needle trajectory are marked. See more examples in the supplementary video.
A tracking demonstration is shown in Fig. 4 including three examples [^1] Additional tracking demonstrations are provided in the supplementary video. that demonstrate different challenges in US needle tracking. The first example of In-plane (static) / 0 / Phantom in the first row shows a scenario with interference from needle traces. MambaXCTrack overcomes the interference from needle traces, which arise from the hollow regions created by needle retraction and have a very similar appearance to the needle tip. This interference has led to STMTrack’s failures, whereas our tracker maintains successful tracking. For procedures done within the tissue, the distraction from noise and artifacts becomes more severe, and the needle visibility is degraded. The second example of In-plane (moving) / 30 / Tissue demonstrates this distraction and degradation, where the background is filled with noise from high-intensity speckles caused by variations in acoustic impedance. The backward procedure of this example also shows the intermittent invisible needle tip, since the imaging plane and needle tend to be misaligned in this dynamic case, where both probe and needle are moving. MambaXCTrack is the only one among the four that maintains successful tracking even when the tip is intermittently invisible. A similar case of tracking the intermittently disappearing needle is also demonstrated in the backward of out-of-plane / 60° / Tissue, where MixFormerV2 lost tracking but MambaXCTrack was able to keep the target tracked.
Our experiments show that MambaXCTrack runs at 34.9 FPS, which meets the real-time requirement for ultrasound needle tracking despite being slower than the MixFormerV2 and SwinTrack. However, our method consistently outperforms these methods in terms of needle tracking accuracy by leveraging pixel-level global modeling with SSM, which efficiently extracts rich semantic cues from neighboring regions. The linear complexity of SSM enables this pixel-level global operation, which can hardly be achieved by a transformer-based method without a significant increase in computational complexity. Therefore, MambaXCTrack achieves an appropriate balance between inference speed and accuracy, as clinical applications demand reliability more over speed gains when real-time thresholds can be met.
@@ III-C Ablation Study
Extensive ablation studies were conducted regarding the baseline model MambaXCTrack against seven variations $v_{1}$ to $v_{7}$, with their configurations and results shown in Tab. III. The first three were adopted to test the proposed SSMX-Corr and CIS. CIS is kept in $v_{1}$ but $m$ is concatenated between $z$ and $x_{i}$, different from the baseline ($[m,z,x_{i}]$), to evaluate different orders of information fusion. CIS is removed in $v_{2}$ and three feature maps are simply concatenated together into a sequence $[m,z,x]$. CIS is also removed in $v_{3}$ and SSMX-Corr is replaced by a convolutional X-Corr [10]. The results of these variations show the effectiveness of CIS and SSMX-Corr. Performance degradation in $v_{1}$ demonstrates that the tracking benefits more from motion information by concatenating $m$ at the beginning of $z$ and $x_{i}$. With CIS removed ($v_{2}$), performance degradation is observed in almost all evaluations, showing that tracking can be improved by building pixel-wise interaction between $z$ and $x$. Performance drop is also observed in $v_{3}$, especially for the tissue experiments, where noise distraction is much more severe. It demonstrates that tracking under challenging environments is greatly improved by learning potential distant visual cues utilizing SSMX-Corr’s long-range modeling capability.
Ablation regarding the proposed low-level motion descriptor was also carried out, namely $v_{4}$ to $v_{7}$. The motion descriptor is removed in $v_{4}$. $v_{5}$ keeps the motion descriptor but uses raw motion without preprocessing it like the baseline. Motion descriptors that cover different time durations are tested in $v_{6}$ (longer time, $T=120$) and $v_{7}$ (shorter time, $T=30$). A considerable degradation is reported in $v_{4}$, demonstrating that tracking is greatly enhanced by learning from historical low-level motion descriptors. Accuracy drops are also reported in almost all metrics of $v_{5}$, showing the improvement brought by extracting low-level descriptors from raw motion. The results of $v_{6}$ and $v_{7}$ demonstrate the length of baseline ($T=60$), which covers 2 seconds, is a more suitable time span for motion descriptors in US needle tracking. This is likely because moderate-duration motion provides sufficient cues while avoiding interference from long-term noise.
@@ III-D Discussion
MambaXCTrack outperforms a series of SOTA trackers including CNN-based trackers and transformer-based trackers. Its superior performance can be attributed to two main factors. Firstly, the SSMX-Corr with CIS effectively addresses the unique domain characteristics of US images, i.e., the local features from the target can become unpredictably unreliable, but searching for semantic cues over long distances significantly enhances tracking accuracy. Studies using needle shaft segmentation for guidance [17, 18] suggest that while the needle tip may occasionally become obscured, the larger needle shaft remains partially visible, providing directional information. Instead of direct segmentation, SSM is employed to endow the model with long-range modeling capabilities, allowing it to learn from distant semantic features including the needle shaft. The existing ConvX-Corr adopted in SiamRPN++ [10] is a local pixel-wise operation that convolves $x$ with kernel $z$. While it is considered effective for open-world tasks, its limited search range proves inadequate for US needle tracking, where images are often degraded by noise and artifacts, rendering local fine features unreliable. With SSMX-Corr, MambaXCTrack overcomes these challenges by identifying long-range semantic cues. By further integrating CIS, SSMX-Corr retains the advantages of ConvX-Corr, gaining positional inductive bias from this convolution-like operation.
@@
The second aspect is the implicit low-level motion descriptor. When the needle temporarily disappears, historical motion becomes more critical than visual information. In such cases, relying on motion information rather than vision is a more effective strategy. Ablation studies demonstrate that integrating historical low-level motion descriptors with an appropriate time duration enhances tracking. SwinTrack [34] also incorporates motion information yet leads to worse performance than ours. This is partly because SwinTrack directly integrates raw motion sequences without calculating relative displacement like MambaXCTrack, which hinders the generalization. Additionally, its feature fusion is less efficient than our SSM, which is inherently suited for long-sequence modeling. STMTrack [33] adopts a template bank to store the historical frames, focusing on historical visual information rather than motion information. However, it performs even worse than SwinTrack [34]. This might be because of error accumulation due to aggregating historical US images, which contain severe noise and artifacts. Yet historical motion suffers less from error accumulation thanks to its simpler distribution and characteristic that does not contain visual features. In challenging environments, leveraging simpler motion information is more effective than integrating visual features. However, while the tracker proposed by Yan et al. [6] also utilizes motion for ultrasound needle tracking, its performance is unsatisfactory. This suboptimal performance is likely attributed to its explicit motion prediction structure, which prevents the model from being fully end-to-end, resulting in error propagation and system instability. Directly predicting future motion can also cause degraded generalizability when encountering unseen videos, which contain motion that can be hard to be adapted by a simple motion predictor. On the contrary, MambaXCTrack leverages implicit motion integration to ensure an end-to-end structure with better generalization.